OpenStudy (anonymous):

rick is riding in a race the equation d=40-1/5 m where d= distance in miles and m= minutes can be used to estimate the number of miles rick has left before reaching the finish line how many minutes has rick been riding when he was 31 miles from the finish line

OpenStudy (anonymous):

Substitute then value distance into your formula and solve for the missing variable

OpenStudy (anonymous):

\[d = 40-\frac{ 1 }{ 5m }\] ?

OpenStudy (anonymous):

thank you im in summer and dont know how to do most of these problems ha

OpenStudy (anonymous):

It is alright, you can learn them in no time. D = distance M = minutes So if you want to find the Minutes at Distance 31 miles then you just substitute 31 for "d" and solve for "m".

OpenStudy (anonymous):

do i - 40 to 30?

OpenStudy (anonymous):

31

OpenStudy (anonymous):

Not sure what you mean?

OpenStudy (anonymous):

Since we want to find the time (m) when d = 31 then we will have to place 31 as "d" into our equation.

OpenStudy (anonymous):

So we now have \[31 = 40 - \frac{ 1 }{ 5m }\] Solve for m

OpenStudy (anonymous):

how do we solve for m ?

OpenStudy (anonymous):

Get m by itself by using algebra. Anything you do to the right side you must do to the left.

OpenStudy (anonymous):

How will you get rid of 40 from the right? And then how will you get rid of -1/5?

OpenStudy (anonymous):

i - the 40 to the 31 and then i dont know how to get rid of the -1/5

OpenStudy (anonymous):

Alright good first start so if we do as you said, \[31 = 40 - \frac{ 1 }{ 5m }\] Subtracting 40 from both sides.\[31-40 = 40 - 40 - \frac{ 1 }{ 5m }\]\[-9 = -\frac{ 1 }{ 5m }\]

OpenStudy (anonymous):

i dont know what to do after that tho

OpenStudy (anonymous):

Now we must get rid of the -1/5 which we can do by multiplying 5 to both sides. \[-5 \times -9 = -5 \times -\frac{ 1 }{ 5m }\]\[-5 \times -9 = \frac{ -5 }{ -5 m }\] Since negatives and negative equals positives and 5 over 5 is 1 we have... \[45 = \frac{ 1 }{ m }\] You must now take the reciprocals of both sides which means flip both sides.

OpenStudy (anonymous):

so it would be 1/m=45

OpenStudy (anonymous):

No, you must take the reciprocals of both which means flipping both the terms.

OpenStudy (anonymous):

i dont understand that part

OpenStudy (anonymous):

\[45 = \frac{ 1 }{ m }\] Which is really \[\frac{ 45 }{ 1 } = \frac{ 1 }{ m }\] Taking the reciprocals of both is flipping both fractions.

OpenStudy (anonymous):

then i cross cancel

OpenStudy (anonymous):

No, flip the fractions first and you will see your result

OpenStudy (anonymous):

so it would be 1/45=m/1

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

Yes that is correct

OpenStudy (anonymous):

ive been on this answer for like 40 min now ha

OpenStudy (anonymous):

\[\frac{ 1 }{ 45 } = \frac{ m }{ 1 }\] Which is the same thing as \[\frac{ 1 }{ 45 } = m\]

OpenStudy (anonymous):

Assuming the original equation and the variable for d are correct

OpenStudy (anonymous):

i just need the answer bro i am like done with this problem please